# SAVOR: Skill Affordance Learning from Visuo-Haptic Perception for Robot-Assisted Bite Acquisition

### v1

@@ Abstract
Robot-assisted feeding requires reliable bite acquisition, a challenging task due to the complex interactions between utensils and food with diverse physical properties. These interactions are further complicated by the temporal variability of food properties—for example, steak becomes firm as it cools even during a meal. To address this, we propose SAVOR, a novel approach for learning skill affordances for bite acquisition—how suitable a manipulation skill (e.g., skewering, scooping) is for a given utensil-food interaction. In our formulation, skill affordances arise from the combination of tool affordances (what a utensil can do) and food affordances (what the food allows). Tool affordances are learned offline through calibration, where different utensils interact with a variety of foods to model their functional capabilities. Food affordances are characterized by physical properties such as softness, moisture, and viscosity, initially inferred through commonsense reasoning using a visually-conditioned language model and then dynamically refined through online multi-modal visuo-haptic perception using SAVOR-Net during interaction. Our method integrates these offline and online estimates to predict skill affordances in real time, enabling the robot to select the most appropriate skill for each food item. Evaluated on 20 single-item foods and 10 in-the-wild meals, our approach improves bite acquisition success by 13% over state-of-the-art (SOTA) category-based methods (e.g. use skewer for fruits). These results highlight the importance of modeling interaction-driven skill affordances for generalizable and effective robot-assisted bite acquisition. Website: https://emprise.cs.cornell.edu/savor/
@@ 2 Related Work
Integrating Vision and Haptics in Robotics. Vision-only manipulation has demonstrated effectiveness in robotic domains such as semantic grasping and deformable object manipulation [16, 17, 9]. However, contact-rich tasks, such as in-hand manipulation [18, 19] and object packing [20], benefit significantly from the integration of vision and haptic information. Previous studies [5, 10, 21, 22, 23, 24, 25, 26, 27, 28] have highlighted the importance of haptic feedback in food manipulation. More recently, Sundaresan et al. [8] and Gordon et al. [10] combine a single image with time-series haptics to learn skewering skills, but rely either on haptics alone or static visual input. In contrast, SAVOR focuses on learning food physical properties from time-series observations of both vision and haptics, allowing us to capture dynamic changes, such as deformation and surface texture variations, which enables a more comprehensive understanding of the food physical properties and their temporal variations.
@@ 3 Problem Formulation
We consider the problem of bite acquisition, where a robot equipped with a utensil needs to acquire food items from a plate in a specified sequence. The user specifies the desired food sequence (e.g., via natural language [3]), and the robot executes one or more actions per item to acquire it. Upon a successful attempt, the item is removed from the tool and the robot proceeds to the next. The objective is to maximize the number of food items acquired within a limited number of attempts.
@@ 4.2 Pre-Deployment: Offline Training SAVOR-Net
We pretrain SAVOR-Net on an existing dataset [43], which contains 400 examples of human skewering various food items, and then fine-tune it on our own dataset, consisting of 300 examples of a robot performing skewering, pushing, twirling, cutting, dipping, and scooping across 20 food items. More details about the model and the training process are available in Appendix B.3
@@ 4.4 During Deployment: Planning
Please select an action from [’skewer, ’scoop’, ’twirl’] to pick up the food item or select
one action from [’cut’, ’dip’, ’push’] to manipulate items to facilitate subsequent
@@
For the low-level skills, we use the skill library from FLAIR [3], which includes four acquisition skills (skewer, twirl, scoop, dip) and two pre-acquisition skills (push, cut), illustrated in Figure 3. Given these low-level skills, SAVOR enables effective skill selection for each food item, despite variations in physical properties and the partial observability present during deployment.
@@ 5 Experiments
Table 1: Quantitative results on bite acquisition. We evaluate our approaches on 10 different dishes (Figure 4) and report success rate.
Plate SAVOR Haptic-only Vision-only FLAIR VLM SayCan End2End
1 10/15 9/17 6/26 9/18 9/19 9/12 9/15
@@ 5.1 Setup
Figure 5: (a) System setup. (b) Average Success Rate using a plastic/metal fork for 10 in-the-wild dishes. (c) SR for 10 in-the-wild dishes. * indicates statistically significant differences ($p<0.05$).
[…]
@@
Figure 7: Generalization performance on seen and unseen food items. We compare SAVOR and SOTA FLAIR across 20 food items. We show SR3 on 10 repetitions per food item. Asterisks (*) indicate statistically significant differences ($p<0.05$).
@@ 5.2 Results and Analysis
Evaluating the Contribution of Visuo-Haptic Perception. We next evaluate how perception modalities affect performance by comparing SAVOR with its vision-only and haptics-only variants (Table 1). SAVOR achieves the highest success rate, confirming that combining vision and haptics improves food property estimation and significantly improves performance from SR1 to SR2 ($p<0.05$). The vision-only variant performs the worst (27.1% average success), struggling with fine-grained utensil–food interactions and often misassigning visually similar items with inaccurate physical properties. For example, strawberries, watermelon, and carrots on Plate 1 appear similar in color, but differ significantly in physical properties. The model frequently overestimates softness for firm items like carrots, leading to repeated skill selection failures. The haptics-only variant performs better (40.1% average SR) but still falls short of SAVOR. The haptics-only variant often fails when food items share similar force signatures, such as tofu and nuts. In these cases, small changes in end-effector pose lead to rapid force increases—either due to skewering through soft items and contacting the plate, or due to deformation of the utensil against firm items. Without visual input, these ambiguities lead to misinterpretation of food properties. By integrating vision and haptics, SAVOR resolves ambiguities, achieving a 51.5% average SR and 87.3% SR3. Figure 5c demonstrates that SAVOR effectively learns from interaction during the initial attempt and adapts its skill selection dynamically, resulting in the greatest improvement in subsequent attempts. These results highlight the critical role of time-series visuo-haptic perception in estimating food affordances.
Evaluating the Overall Framework. To evaluate our full framework, we compare SAVOR with the SOTA method (FLAIR), a vision-only variant using a VLM, SayCan and the end-to-end method. SAVOR achieves the highest average SR (51.5%) and 87.3% SR3 across 10 diverse plates (Table 1).
FLAIR selects skills based on predefined food categories, achieving 38.2% average SR and 73.4% SR3. It performs well on stereotypical items (e.g., banana), but fails on atypical foods such as raw avocado, where category-based reasoning leads to inappropriate actions (e.g., skewering instead of scooping). VLM improves upon FLAIR, achieving 44.1% average SR and 81.0% SR3 by inferring food properties from visual input. However, it struggles with visually ambiguous or dynamically changing items, such as misclassifying avocado or failing to detect firmness changes in cooling chicken nuggets. Similarly, SayCan achieves only 36.2% average SR and 73.4% SR3, often oscillating between skewer and scoop skills. Without compact physical property representations, the end-to-end policy generalizes poorly, reaching 42.7% average SR and 81.8% SR3.
@@ A.2 Ablation Study on Tool Calibration
We provide detailed results of the ablation study on the calibration process for the 10 in-the-wild dishes (Table 3). Compared to the uncalibrated baseline, tool calibration significantly improves performance for both the plastic and metal forks. Specifically, with calibration, the plastic fork (PF) achieves a 51.5% average success rate and 87.3% SR3, compared to only 38.7% and 75.9% without calibration (PF-wo). Similarly, the metal fork (MF) benefits from calibration, improving from 83.5% SR3 to 93.7%. These results demonstrate that understanding tool capabilities through calibration helps the planner avoid infeasible actions, such as skewering tofu with a metal fork (Plate 8) or skewering firm steak with a plastic fork (Plate 10), thereby improving skill selection and acquisition success.
@@
Plate Type Methods
Plate Items Visual Haptic PF PF-wo MF MF-wo
1 strawberries*, watermelon, carrots Similar Diverse 10/15 9/19 11/15 11/15
2 tomatoes*, broccoli, carrots Similar Diverse 7/13 4/17 6/15 5/17
3 avocado*, banana, sauce Diverse Diverse 7/11 5/15 7/10 7/9
4 muffin*, cake*, jello Diverse Diverse 6/13 6/13 7/11 7/11
5 cookie*, bread*, cheese Diverse Diverse 6/13 7/12 7/12 7/12
6 roasted turkey, chicken nuggets*, mashed potatoes, green beans* Diverse Diverse 10/18 8/21 10/17 10/18
7 salmon, mushrooms Diverse Similar 5/17 6/17 7/17 7/15
8 chicken breast, tofu, mushrooms Diverse Diverse 5/9 3/12 6/9 5/11
9 chicken, broccoli*, noodles* Diverse Diverse 7/10 7/10 6/12 6/12
10 steak*, broccoli*, mashed potatoes Diverse Diverse 6/16 5/19 8/13 7/13
Average Success Rate (%) 51.5% 38.7% 56.1% 54.1%
@@ A.3 Crumbly and Soft Foods
We evaluate our approach on foods that are particularly soft or crumbly, such as tofu and crackers. For crackers, our system selects scooping in 80% of trials and attempts skewering in 20% in the initial attempt. After a skewering attempt, SAVOR identifies low softness, logs the failure in the attempt history, and switches to scooping. For tofu, we test variants from extra soft to super firm. The system initially selects skewering for all tofu types but switches to scooping for extra soft tofu. We achieve a 70% average SR for crackers and a 95% average SR for tofu. These results suggest that our method can adapt to diverse food items when informed by interaction feedback.
@@
A.4 Open-loop Nature of Skill Execution
@@
A.5 Multi-food Interactions on Cluttered Plates
@@ B.3.1 Model Architecture
SAVOR-Net uses separate encoders for each of the time series and further splits the RGB-D inputs into RGB and depth for separate encoding. The encoder for RGB images is a pre-trained ResNet50 followed by a two-layer MLP. The encoder for depth images is a 4-layer convolutional neural network, followed by a two-layer MLP, where each convolutional layer has a 3$\times$3 kernel and is followed by Leaky ReLU activation. The encoder for haptics $F_{t}$ is a two-layer MLP and the encoder for end-effector poses $P_{t}$ is a two-layer MLP. Each encoder outputs a vector in $\mathbb{R}^{128}$. The four vectors are concatenated into a unified multimodal representation and then passed to an LSTM with 2 layers and a hidden size of 512. A three-layer MLP takes output from the LSTM and produces the final output $\psi_{t}$.
@@ B.3.2 Training
SAVOR-Net is trained using cross-entropy loss with the hyperparameters listed in Table 4. Training is conducted on an NVIDIA RTX 3070 GPU and completes in approximately 50 minutes.
@@ B.4.4 Skill Selection
Please select an action from [’skewer, scoop’, twirl’, dip’] to pick up the food item. Notice that if all acquisition skills are not immediately feasible, please select an action from [’cut’, push’] to rearrange or manipulate items to facilitate subsequent acquisition. Always follow the format: Reasoning: <your reason>. Answer: <your answer>.
@@
Please select an action from [’skewer, scoop’, twirl’, dip’] to pick up the food item. Notice that if all acquisition skills are not immediately feasible, please select an action from [’cut’, push’] to rearrange or manipulate items to facilitate subsequent acquisition. Always follow the format: Reasoning: <your reason>. Answer: <your answer>.
@@
Please select an action from [’skewer, scoop’, twirl’, dip’] to pick up the food item. Notice that if all acquisition skills are not immediately feasible, please select an action from [’cut’, push’] to rearrange or manipulate items to facilitate subsequent acquisition. Always follow the format: Reasoning: <your reason>. Answer: <your answer>.

### v2

@@ Abstract
Robot-assisted feeding requires reliable bite acquisition, a challenging task due to the complex interactions between utensils and food with diverse physical properties. These interactions are further complicated by the temporal variability of food properties—for example, steak becomes firm as it cools even during a meal. To address this, we propose SAVOR, a novel approach for learning skill affordances for bite acquisition—how suitable a manipulation skill (e.g., skewering, scooping) is for a given utensil-food interaction. In our formulation, skill affordances arise from the combination of tool affordances (what a utensil can do) and food affordances (what the food allows). Tool affordances are learned offline through calibration, where different utensils interact with a variety of foods to model their functional capabilities. Food affordances are characterized by physical properties such as softness, moisture, and viscosity, initially inferred through commonsense reasoning using a visually-conditioned language model and then dynamically refined through online visuo-haptic perception using SAVOR-Net during interaction. Our method integrates these offline and online estimates to predict skill affordances in real time, enabling the robot to select the most appropriate skill for each food item. Evaluated on 20 single food items and 10 in-the-wild meals, our approach improves bite acquisition success rate by 13% over state-of-the-art (SOTA) category-based methods (e.g. use skewer for fruits). These results highlight the importance of modeling interaction-driven skill affordances for generalizable and effective robot-assisted bite acquisition. Website: https://emprise.cs.cornell.edu/savor
@@ 2 Related Work
Integrating Vision and Haptics in Robotics. Vision-only manipulation has demonstrated effectiveness in robotic domains such as semantic grasping and deformable object manipulation [16, 17, 9, 18, 19]. However, contact-rich tasks, such as in-hand manipulation [20, 21] and object packing [22], benefit significantly from the integration of vision and haptic information. Previous studies [5, 10, 23, 24, 25, 26, 27, 28, 29, 30, 31] have highlighted the importance of haptic feedback in food manipulation. More recently, Sundaresan et al. [8] and Gordon et al. [10] combine a single image with time-series haptics to learn skewering skills, but rely either on haptics alone or static visual input. In contrast, SAVOR focuses on learning food physical properties from time-series observations of both vision and haptics, allowing us to capture dynamic changes, such as deformation and surface texture variations, which enables a more comprehensive understanding of the food physical properties and their temporal variations.
@@ 3 Problem Formulation
We consider the problem of bite acquisition, where a robot equipped with a utensil needs to acquire food items from a plate in a specified sequence. The user specifies the desired food sequence (e.g., via natural language [3]), and the robot executes one or more actions per item to acquire it. Upon a successful attempt, the item is removed from the tool, and the robot proceeds to the next. The objective is to maximize the number of food items acquired within a limited number of attempts.
@@ 4.2 Pre-Deployment: Offline Training SAVOR-Net
We pretrain SAVOR-Net on an existing dataset [47], which contains 400 examples of human skewering various food items, and then fine-tune it on our own dataset, consisting of 300 examples of a robot performing skewering, pushing, twirling, cutting, dipping, and scooping across 20 food items. More details about the model and the training process are available in Appendix B.3.
@@ 4.4 During Deployment: Planning
Please select an action from [’skewer, ’scoop’, ’twirl’, ’dip’] to pick up the food item or
select one action from [’cut’, ’push’] to manipulate items to facilitate subsequent
@@
For the low-level skills, we use the skill library from FLAIR [3], which includes four acquisition skills (skewer, twirl, scoop, dip) and two pre-acquisition skills (cut, push), illustrated in Figure 3. Given these low-level skills, SAVOR enables effective skill selection for each food item, despite variations in physical properties and the partial observability present during deployment.
@@ 5 Experiments
Table 1: Quantitative results on bite acquisition. We evaluate our approach on 10 different dishes (Figure 4) and report both Attempt Efficiency (AE) and Success Rate (SR), averaged across plates.
Plate #Items Acquired / #Total Attempts
SAVOR Haptic-only Vision-only FLAIR VLM SayCan End2End
1 10/15 9/17 6/26 9/18 9/19 9/22 9/15
@@ 5.1 Setup
Figure 5: (a) System setup. (b) SR3 using a plastic/metal fork for 10 in-the-wild dishes. (c) SR{1,2,3} for 10 dishes. * indicates statistically significant differences ($p<0.05$).
Evaluation Metrics. We evaluate each method using Success Rate. A trial is considered successful if the utensil picks up the food item and retains it for at least 5 seconds. Failed attempts leave the item on the plate, allowing up to three re-attempts before it is manually removed and recorded as a failure. To capture the efficiency of acquisition, we define $\text{Average Attempt Efficiency (average AE)}=\frac{\#\text{Items Acquired}}{\#\text{Total Attempts}}$ to quantify the proportion of successful acquisitions relative to total attempts across all plates. To further quantify overall performance, we define SR1 as the proportion of items acquired within the first attempt in a meal, i.e., SR1 = $\frac{\#\text{Items Acquired}}{\#\text{Total Items}}$ within one attempt. Similarly, SR2 and SR3 represent the proportion of items acquired within two and three attempts, respectively.
@@
Figure 7: Generalization performance on seen and unseen food items. We compare SAVOR and SOTA FLAIR across 20 food items. We show SR3 on 10 episodes per food item, where each episode allows up to 3 attempts. Asterisks (*) indicate statistically significant differences ($p<0.05$).
@@ 5.2 Results and Analysis
Evaluating the Contribution of Visuo-Haptic Perception. We next evaluate how perception modalities affect performance by comparing SAVOR with its vision-only and haptics-only variants (Table 1). SAVOR achieves the highest success rate, confirming that combining vision and haptics improves food property estimation and significantly improves performance from SR1 to SR2 in Figure 5c ($p<0.05$). The vision-only variant performs the worst (27.1% average AE), struggling with fine-grained utensil–food interactions and often misassigning visually similar items with inaccurate physical properties. For example, strawberries, watermelon, and carrots on Plate 1 appear similar in color, but differ significantly in physical properties. The model frequently overestimates softness for firm items like carrots, leading to repeated skill selection failures. The haptics-only variant performs better (40.1% average AE) but still falls short of SAVOR. The haptics-only variant often fails when food items share similar force signatures, such as tofu and nuts. In these cases, small changes in end-effector pose lead to rapid force increases—either due to skewering through soft items and contacting the plate, or due to deformation of the utensil against firm items. Without visual input, these ambiguities lead to misinterpretation of food properties. By integrating vision and haptics, SAVOR resolves ambiguities, achieving a 51.5% average AE and 87.3% SR3. Figure 5c demonstrates that SAVOR effectively learns from interaction during the initial attempt and adapts its skill selection dynamically, resulting in the greatest improvement in subsequent attempts. These results highlight the critical role of time-series visuo-haptic perception in estimating food affordances.
Evaluating the Overall Framework. To evaluate our full framework, we compare SAVOR with the SOTA method (FLAIR), a vision-only variant using a VLM, SayCan and the end2end method. SAVOR achieves the highest average AE (51.5%) and 87.3% SR3 across 10 diverse plates (Table 1).
FLAIR selects skills based on predefined food categories, achieving 38.2% average AE and 73.4% SR3. It performs well on stereotypical items (e.g., banana), but fails on atypical foods such as raw avocado, where category-based reasoning leads to inappropriate actions (e.g., skewering instead of scooping). VLM improves upon FLAIR, achieving 43.8% average AE and 81.0% SR3 by inferring food properties from visual input. However, it struggles with visually ambiguous or dynamically changing items, such as misclassifying avocado or failing to detect firmness changes in cooling chicken nuggets. Similarly, SayCan achieves only 36.2% average AE and 73.4% SR3, often oscillating between skewer and scoop skills. Without compact physical property representations, the end-to-end policy generalizes poorly, reaching 42.7% average AE and 81.8% SR3.
@@ A.2 Ablation Study on Tool Calibration
We provide detailed results of the ablation study on the calibration process for the 10 in-the-wild dishes (Table 3). Compared to the uncalibrated baseline, tool calibration significantly improves performance for both the plastic and metal forks. Specifically, with calibration, the plastic fork (PF) achieves a 51.5% average attempt efficiency and 87.3% SR3, compared to only 38.7% and 75.9% without calibration (PF-wo). Similarly, the metal fork (MF) benefits from calibration, improving from 83.5% SR3 to 93.7%. These results demonstrate that understanding tool capabilities through calibration helps the planner avoid infeasible actions, such as skewering tofu with a metal fork (Plate 8) or skewering firm steak with a plastic fork (Plate 10), thereby improving skill selection and acquisition success.
@@
Plate #Items Acquired / #Total Attempts
PF PF-wo MF MF-wo
1 10/15 9/19 11/15 11/15
2 7/13 4/17 6/15 5/17
3 7/11 5/15 7/10 7/9
4 6/13 6/13 7/11 7/11
5 6/13 7/12 7/12 7/12
6 10/18 8/21 10/17 10/18
7 5/17 6/17 7/17 7/15
8 5/9 3/12 6/9 5/11
9 7/10 7/10 6/12 6/12
10 6/16 5/19 8/13 7/13
Average AE (%) 51.5% 38.7% 56.1% 54.1%
@@ A.3 Crumbly and Soft Foods
We evaluate our approach on foods that are particularly soft or crumbly, such as tofu and crackers. For crackers, our system selects scooping in 80% of trials and attempts skewering in 20% in the initial attempt. After a skewering attempt, SAVOR identifies low softness, logs the failure in the attempt history, and switches to scooping. For tofu, we test variants from extra soft to super firm. The system initially selects skewering for all tofu types but switches to scooping for extra soft tofu. We achieve a 70% average attempt efficiency for crackers and a 95% average attempt efficiency for tofu. These results suggest that our method can adapt to diverse food items when informed by interaction feedback.
@@
A.4 VLMs Failures
We use a single VLM (GPT-4V) for both state estimation and skill planning, instantiated separately as GPT-Property and GPT-Skill. While the VLM itself is static, we augment its context after each interaction with the latest observations. After every interaction, GPT-Property can update food properties. GPT-Skill can also adjust skill choice because some of its inputs are from GPT-Property.
[…]
A.5 Open-loop Nature of Skill Execution
@@
A.6 Multi-food Interactions on Cluttered Plates
@@ B.3.1 Model Architecture
SAVOR-Net uses separate encoders for each of the time series and further splits the RGB-D inputs into RGB and depth for separate encoding. The encoder for RGB images is a pre-trained ResNet50 followed by a two-layer MLP. The encoder for depth images is a 4-layer convolutional neural network, followed by a two-layer MLP, where each convolutional layer has a $\times3$ kernel and is followed by Leaky ReLU activation. The encoder for haptics $F_{t}$ is a two-layer MLP and the encoder for end-effector poses $P_{t}$ is a two-layer MLP. Each encoder outputs a vector in $\mathbb{R}^{128}$. The four vectors are concatenated into a unified multimodal representation and then passed to an LSTM with 2 layers and a hidden size of 512. A three-layer MLP takes output from the LSTM and produces the final output $\psi_{t}$.
@@ B.3.2 Training
SAVOR-Net($\sim$14M parameters) is trained using cross-entropy loss with the hyperparameters listed in Table 4. Training is conducted on an NVIDIA RTX 4090 GPU and completes in approximately 40 minutes.
@@ B.4.4 Skill Selection
Please select an action from [’skewer, scoop’, twirl’, dip’] to pick up the food item or
select one action from [’cut’, push’] to manipulate items to facilitate subsequent
@@
Please select an action from [’skewer, scoop’, twirl’, dip’] to pick up the food item or
select one action from [’cut’, push’] to manipulate items to facilitate subsequent
@@
Please select an action from [’skewer, scoop’, twirl’, dip’] to pick up the food item or
select one action from [’cut’, push’] to manipulate items to facilitate subsequent